OpenStudy (anonymous):

Abel makes the following statement: "I have a 30% chance of being selected as the captain of the baseball team and a 20% chance of being elected student body president." What is the probability that Abel will be selected as the captain of the baseball team and will be elected student body president? 6% 10% 20% 50%

OpenStudy (anonymous):

I will medal and fan!!! Please help asap :)

OpenStudy (faiqraees):

P(Captain and President) = P(Captain) * P(President) = 0.3 * 0.2 = ?

OpenStudy (faiqraees):

No

OpenStudy (faiqraees):

* means multiply

OpenStudy (anonymous):

oh right i wasn't paying attention -_- so it would be 6%??

OpenStudy (faiqraees):

Yes

OpenStudy (anonymous):

thank you!! are you positive:)?

OpenStudy (faiqraees):

What do you need help for?

OpenStudy (anonymous):

nothing really anymore you already helped:) but @DARTHVADER2900 has been watching this question for like 20 minutes now lol